OpenStudy (anonymous):

What steps do I take to solve this problem that I can use with any substitution problem? Solve using substitution. -x + 3y = -4 y = -1 (, )

OpenStudy (anonymous):

I like your profc it reminds me of someone else!!!!! -x+ 3(-1)=-4 -x-3=-4 +3 +3 x=-1 I think that is correct

OpenStudy (anonymous):

thx